QUESTION 4 Find the equivalent capacitance for the circuit shown. 6 F 3 uF 7 1 uF Ceg- UF
Answer #1

: 6uF leq But I LUF Parallel Cu = C,+C₂ = 3+1 Cu =4uF buf Ceq Series Су 4UF Caly Ceq = 6x4 Cutly 4+6 Ceq=2.4MF ich is C₂ para